OpenStudy (anonymous):

Write an expression that is equivalent to three over five(3y + 15). three over fivey + 9 three over five y + 15 nine over fivey + 9 nine over fivey + 15

OpenStudy (anonymous):

@133

OpenStudy (anonymous):

@ayyyyyyyyy25

OpenStudy (anonymous):

@hlee0919

OpenStudy (texaschic101):

3/5 (3y + 15) what we need to do is distribute the 3/5 through the parenthesis. That just means multiply everything in the parenthesis by 3/5. 3/5 * 3y = 9/5y 3/5 * 15 = 9 so we have : 9/5y + 9

OpenStudy (anonymous):

ooohhhhh that is why well thank u soo much

OpenStudy (texaschic101):

sure thing :)